Find a number c that satisfies the Mean Value Theorem for the function
F(x) = x ^3 -2x on the interval [1,2]

Round your answers to 4 decimal places
1
Expert's answer
2011-03-10T10:03:10-0500
According to the Mean Value Theorem,
f'(c) = (f(b) - f(a))/(b-a) = (f(2) - f(1))/(2-1) = 4 +1 = 5
f'(c) = 3c2 - 2 = 5
3c2 = 7
c = √(7/3) = +1.5275